A shop has determined that the number of bottles of water they sell on a summer day is a function of the day's high temperature, and is approximately given by N(H)=2H+31, where H is the day's high temperature, and 80≤H≤110. Suppose that the predicted high temperatures for the net 7 days are given in the table below.

t (day) High temp, H (degrees F) 1 95 2 96 3 97 4 98 5 99 6 100 7 101

Q2.1 Part a)

Express the high temperature as a function of the day.

Q2.2 Part b)

Use this function and the function provided to create a function that represents the number of bottles of water sold as a function of the day.

Q2.3 Part c)

Pick one ordered pair on the graph of the new function, and explain what it tells you in practical terms.
Answer #1

ay et Day Temperature - H *, H = tz H₂ = 1,95 2,96 H-H, - H2-H, (tot.) tz-t, H-95 = 96-95 (tall) 2-1 H-953 t- H= t +44 b ) .